OpenStudy (anonymous):

A cheetah runs along the path such that the displacement vs. time (meters vs. seconds) graph is given by the equation: x(t)=(6/π)sin-1t•cos-1t. What is the instantaneous velocity of the cheetah at the point (1/2, π/3)?

OpenStudy (mos1635):

x(t)=(6/π)sin-1t•cos-1t x(t)=(6/π)sin(-t)•cos(-t) ?????????????

OpenStudy (mos1635):

is that the equation??

OpenStudy (mos1635):

x(t)=(6/π)sin(-t)•cos(-t)